# Augmenting the action space with conventions to improve multi-agent cooperation in Hanabi

### v2

@@ A.2 Hanabi
The principles for Hanabi are based on those defined by the H-Group. We strongly encourage the reader to read their documentation, specifically chapters I – VIII, on Hanabi, conventions, and convention principles for an in-depth discussion (with examples) for each included principle and their corresponding conventions discussed below. As discussed in Section 4.3, the implemented conventions only consist of the most fundamental principles, since these are generally considered standardised within the Hanabi community. Majority of the listed conventions are single-or two-step conventions, however the Prompt and the Finesse are multistep. Additionally, the H-Group discuss more advance multistep conventions useful for future considerations. Note that the prompt and finesse is only applicable in three-to-five players Hanabi, however there does exist a self-prompt for two-player Hanabi which reduces the convention to a two-step convention, and is included in our implementation.
@@ Appendix C Statistical Significance Test
These results demonstrate that majority of the Rainbow with conventions agents’ performances are statistically different from existing research conducted on the Hanabi problem for self-play as well as cross-play agents. The Rainbow with conventions agent’s scores are statistically similar to that of baseline Rainbow two players and ACHA three players, however conventions allow the Rainbow algorithm to reach these scores significantly faster (5x compared to baseline Rainbow and 1000x compared to ACHA). Furthermore, even though these two instances are statistically similar, the agents learn completely different policies, as seen by the cross-play performance and its statistical differences.

### v3

@@ A.2 Hanabi
The principles for Hanabi are based on those defined by the H-Group. We strongly encourage the reader to read their documentation, specifically chapters I – VIII, on Hanabi, conventions, and convention principles for an in-depth discussion (with examples) for each included principle and their corresponding conventions discussed below. As discussed in Section 4.3, the implemented conventions only consist of the most fundamental principles, since these are generally considered standardised within the Hanabi community. Majority of the listed conventions are single-or two-step conventions, however the Prompt and the Finesse are multistep. Additionally, the H-Group discuss more advance multistep conventions useful for future considerations. Note that the prompt and finesse is only applicable in three-to five-player Hanabi, however there does exist a self-prompt for two-player Hanabi which reduces the convention to a two-step convention, and is included in our implementation.
@@ Appendix C Statistical Significance Test
These results demonstrate that majority of the Rainbow with conventions agents’ performances are statistically different from existing research conducted on the Hanabi problem for self-play as well as cross-play agents. The Rainbow with conventions agent’s scores are statistically similar to that of baseline Rainbow two-players and ACHA three-players, however conventions allow the Rainbow algorithm to reach these scores significantly faster (5x compared to baseline Rainbow and 1000x compared to ACHA). Furthermore, even though these two instances are statistically similar, the agents learn completely different policies, as seen by the cross-play performance and its statistical differences.